OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

ah, first we find the slope of the line \[4x - 3y = -12\] well, lets get it to the form of \[y = mx + c \] so \[4x - 3y = -12\] add 3y to both sides to get \[4x - 3y + 3y = -12 + 3y\] which is the same as \[4x = 3y -12\] and now add 12 to both sides so \[4x +12 = 3y -12 + 12\] or \[4x + 12 = 3y\] dividing both sides by three, \[4x/3 + 12/3 = 3y / 3 \] or \[y = 4/3 + 4\] so we know that m = 4/3 now you have a line whose slope is 4/3 and it passes through (-3, 4). How do you write an equation for that?